# beSnake: A routing algorithm for scalable spin-qubit architectures

### v1

@@ V Simulations and Evaluations
With the following simulations, our goal is to test beSnake’s different functionalities and create valuable performance insights under various use cases. More specifically, we test beSnake’s capabilities with different levels of shortest path optimizations, SWAP replacement options, as well as its behavior on more connected topologies and ranges of qubit densities. Based on these insights, we move on to thoroughly compare it with the SBS algorithm on synthetic and real quantum circuits with up to 1000 qubits.
@@ V.1 Experimental Setup
The beSnake and SBS routing algorithms have been incorporated into the SpinQ compilation framework [1] and executed single-threaded on a 2.4GHz and 750GB memory server running Python 3.6.8.
It should be noted that routing times can vary depending on the server’s total load and dynamic job scheduling at the time of submission. We have taken measures to maintain good time consistency between our simulations by submitting jobs in dedicated resources with no influence from other jobs.
As for overhead metrics, we have used the same gate and depth overhead definitions as in [1]. Gate overhead is calculated as the percentage relation of additional gates incorporated by the router divided by the number of gates (after decomposition). Depth overhead is expressed as the percentage of extra depth generated by the router in relation to the initial ideal circuit depth (after decomposition).
@@
Both algorithms are benchmarked with randomly generated circuits as well as eleven well-known quantum algorithms. Regarding the randomly generated quantum circuits, they consist of two-qubit and shuttle-based Z gates in three different ratios (25, 50, and 75 two-qubit gate percent), and are sampled ten times for each data point. With this choice of random algorithms, which contain many combinations of the two gate types parallelized in different ways, we aim to stress-test both routing algorithms. As for the number of gates, we fixed it at 3000 gates since its impact on the gate and depth overhead is negligible, according to [1]. The eleven real quantum algorithms were scaled up to 1000 qubits to observe their performance in Section V.3.2. These algorithms were taken from the Qlib [51], Revlib [52], MQT Bench [53] and qbench [54] libraries.
@@ V.2 Benchmarking beSnake
This section focuses on testing beSnake’s various functionalities. Specifically, we used four time limits dictating the shortest path optimization level and compared them with taking only one shortest path in Section V.2.1. In Section V.2.2, we test the behavior of the SWAP replacement option and observe the scaling on four SWAP fidelities while maintaining the shuttle fidelity constant. Then, in Section V.2.3, we are investigating beSnake’s time performance on more connected topologies, while in Section V.2.4, we are fixing the topology sizes but vary the qubit density.
@@ V.2.1 Adjusting the shortest path optimization
In the following simulations, we explore the reduction in gate overhead under various time limits of the shortest path optimization, as outlined in Section IV.5. Therefore, in Figure 7, we compared the gate overhead (in blue) and total routing time (in red) for , , , and seconds time limits imposed to Networkx.all_shortest_paths() function against just taking one shortest path. The random circuits’ gates have been fully (ideally) parallelized based only on their dependences, and beSnake is configured without the ability of SWAP replacements (see Section IV.5), such that no other factor can influence the results. The underlying question to answer with this simulation is:
In Figures 6(a), 6(b) and 6(c), we can observe the gate overhead for 25, 50, and 75 two-qubit gate percentages, respectively. We observe that in all cases, the increased time limits yield a benefit in terms of gate overhead compared to taking only one shortest path, but the benefits are negligent with higher time limits. As expected, higher two-qubit gate percentages result in higher gate overhead and routing time. Thus, the gate overhead gap between the one shortest path and the time limit simulations is more prominent in higher percentages. However, the overall routing time is severely increasing for each time limit. Finding only one shortest path appears polynomial in time over the number of qubits with a small relative gate overhead reduction of , , and on average for 25, 50, and 75 two-qubit gate percentage, respectively. To solidify this point, in Figure 6(d), we isolated the one shortest path option, clearly showing a nearly linear behavior. A 0.05-second time limit is thus sufficient; increasing it beyond this limit does not yield significant gate overhead benefits, and further increases negative impacts on the total routing time.
@@ V.2.2 SWAP replacement
In this simulation, the behavior of the SWAP replacement option on all three metrics is investigated. To test the effect of different fidelity scales, we have fixed the shuttle to , but tested SWAP gates with , , , and fidelities. As explained in Section IV.5, the SWAP replacement is triggered when the accumulated shuttle fidelity is lower than the SWAP fidelity. Intuitively, we expect replacements to be triggered more often with higher SWAP fidelities. Finally, the gates in the random circuits are serialized in order to pronounce the effect in both overheads (when replacing multiple shuttles at once), and we fixed a 0.05-second time limit for the shortest path optimization. The underlying question to answer with this simulation is:
In Figure 8, SWAP replacements improved all overheads for every two-qubit gate percentage. As expected, the SWAP replacement was utilized more often with higher SWAP fidelities and higher two-qubit gate percentages. However, there is no significant overhead difference shown in the figures for higher than (see the clustered lines), potentially due to the maximum accumulated shuttle fidelity being somewhere between and percent. In particular, three shuttles are needed to obtain an accumulated fidelity between and . This indicates that the BFS exploration never reached more than three layers, equivalent to three consecutive shuttles (similar to the example in Figure 3(b)). Unexpectedly, we observe that the routing time has not worsened but mostly improved for higher than fidelity, possibly due to less time taken to update the intermediate representation data structure of the circuit with just one gate (a SWAP) instead of multiple shuttles. Therefore, the SWAP replacement functionality can be used without an execution time penalty in most cases. It should be noted that, given a higher qubit density on the topology, there can be more consecutive shuttles at one time, which will lower the requirement for such high SWAP fidelities in order to trigger the replacement.
@@ V.2.3 Behavior on more connected topologies
In the following experiments, we are interested in seeing how beSnake performs on topologies with higher connectivity. The comparison is made between a squared grid topology (i.e., each qubit is coupled with four neighboring qubits, and with two at the corners) and a squared grid topology in which diagonal edges in every direction have been added (i.e., each qubit is coupled with eight neighboring qubits, and with three at the corners). Of course, it is expected that there will be an improvement in both overheads on the topologies with more connections. However, the increased graph size and solution space might negatively impact beSnake’s speed in exploring solutions. The random circuits are given the same way as in Section V.2.1 as well as the configuration of beSnake with a 0.05 second time limit. The underlying question to answer with this simulation is:
As can clearly be seen in Figure 9, beSnake can obtain significantly better gate and depth overhead as it was expected, and it can also achieve this in less time. In particular, beSnake is, on average , , and faster for 25, 50, and 75 two-qubit gate percentages, respectively. The benefit is more pronounced with higher two-qubit gate percentages, which is expected. The faster times are attributed to the shorter traversed paths, hence fewer steps to take and fewer BFS exploration layers.
@@ V.2.4 Qubit density
Unlike the other simulations using a checkerboard pattern physical initial placement, we will try a different placement to vary the qubit densities in these. The topology sizes are fixed, and qubits are placed next to each other in a left-to-right bottom-up approach until a particular density is reached. As for the executed circuits and beSnake’s configuration, they are similar to Section V.2.1 with a 0.05 second time limit. The underlying question to answer with this simulation is:
In Figure 10, we have simulated seven levels of densities for four different topology sizes: a 25, 49, 64, and 100 site topology. On the y axes, we can observe the normalized mean values between 0.1 and 1 for all three metrics. Focusing on the 25-site topology, routing time (in green) increases exponentially after density, which is expected primarily due to the additional time needed to finish each BFS exploration. This becomes more severe at density, where there are only empty locations on the topology. Observing the routing time for the larger topology sizes, we see a steeper upward trend after . However, the routing time behavior until appears linear for all sizes. Notably, it becomes flatter for larger sizes, indicating that beSnake performs better on larger topologies. This shows that beSnake’s qubit density sweet spot is around .
It should be noted, however, that for another experimental setup in regards to the hardware used to run beSnake, we might obtain a different trend and possibly an improved one if beSnake becomes multithreaded, for instance. Additionally, changing device characteristics can change beSnake’s performance. For example, increasing the connectivity will favor the run time based on the simulations in Section V.2.3. Having said that, the total run time might be reduced equivalently for all densities, and therefore, the same behavior at might still appear. Whether this sweet spot (or around that region) is universally found for any given setup is left open for future work.
Focusing now on both overheads, we overall observe that beSnake favors gate overhead (in blue) over depth overhead (in orange), which is a direct outcome of the core strategy described in Section IV. Then, depth overhead follows a linear trend for all sizes. Surprisingly, gate overhead creates a slight curve gravitating towards qubit density. Therefore, this reinforces the previous conclusion that beSnake can operate well for more than qubit density and particularly for our particular setup. Finally, the simulations show that in terms of gate and depth circuit overhead, there are no exponential penalties for higher than densities.
@@ V.3 Comparison between beSnake and SBS
In this final section, we will compare the performance of both routers for random and real circuits for up to 1000 qubits. To facilitate a fair comparison (named later as ”beSnake fair”), the circuits are given with their gates as parallelized as possible based on SBS’s maximum capability [1]. The SWAP replacement of beSnake has been deactivated as well because SBS does not support it. Additionally, beSnake’s shortest-path optimizations are disabled, and we only use one shortest path (see Section V.2.1). As noted before, beSnake can handle more complex routing tasks with ideally scheduled circuits based purely on their gate’s dependencies, and, for comparison purposes, we will provide such results in Section V.3.2, named as ”beSnake full”.
@@ V.3.1 Randomly generated circuits with up to 1000 qubits
In Figures 10(a) and 10(b), the gate overhead and depth overhead of both algorithms (marked SBS and beSnake) are shown for the different two-qubit gate percentages (in yellow, blue, pink). We can observe beSnake (circle markers) a relative improvement in gate overhead of and in depth overhead of on average compared to SBS (square markers). Additionally, we observe beSnake’s relative benefit in both overheads increasing with higher qubit counts. Looking at the three two-qubit gate percentages, we see an increased gate overhead improvement with higher percentages (see the distance between curves of the same color). More specifically, the average relative improvement is calculated at % for 25%, for 50%, and for 75%, and for depth overhead is calculated at for 25%, for 50%, and % for 75%. beSnake will only be, on average, slower than SBS at 25%, at 50%, and at 75%. Notably, routing time starts converging around and after 900 qubits, indicating that beSnake takes almost the same amount of time at 1000 qubits, especially for more complex routing tasks (i.e., 75 two-qubit gate percentage shown in Figure 10(c)). Considering the performance gains, beSnake is a more attractive solution for large-scale architectures overall. Finally, as discussed before, in these comparison simulations, beSnake is not utilized fully; hence, we expect even better performance for more parallelized circuits.
@@ V.3.2 Real algorithms with up to 1000 qubits
In the following simulations, we compare the performance of SBS over beSnake for real algorithms with up to 1000 qubits. Previously, the random set of circuits was produced in order to stress the capabilities of both routing algorithms; however, the results might not be representative of routing for real algorithms, which might exhibit different circuit characteristics. Similarly to Section V.3.1, we will compare both routers fairly to see their relative performance. However, we will also assume ideally parallelized circuits and enable the SWAP replacement option (at 99.95% SWAP fidelity) in order to test beSnake at its full capacity. The underlying question to answer with this simulation is:
@@
Table 2 summarizes the relative performance of beSnake over SBS (used as a baseline) in two configurations, one fair comparison, ”beSnake fair”, and one scenario with beSnake’s full capacity, ”beSnake full”. Due to algorithm availability, we have specified each algorithm’s qubit range and incremented qubit steps on the second column. However, it should be noted that the number of gates, whether two-qubit gates or other types, are not the only performance indicators or predictors. To get a full picture of their features, a deeper analysis has to be conducted with the help of the algorithm’s internal interaction and dependency graph [1, 54]. In this simulation, we are focusing on the performance comparison of the two routers rather than on specific quantum algorithm performance or how these algorithms scale.
@@
Moving on to the rest of the columns, we provide the relative performance of beSnake over SBS for gate overhead, depth overhead, and routing time as rGO, rDO, and rTime, respectively. For example, beSnake obtained a 16% gate overhead and 29% depth overhead improvement over SBS under a fair comparison for Grover’s algorithm. As for the rTime, beSnake is 42% slower than SBS. Looking only at the fair comparisons overall, we observe an improvement in both rGO and rDO across all algorithms, with up to a 67% improvement in rGO and up to 45% improvement in rDO. On the one hand, the routing time results of Deutsch-Josza (up to 1000 qubits) and Bernstein-Vazirani (up to 50 qubits) seem relatively high for their low average two-qubit and Z gate values. Evidently, these algorithms do not require a lot of routing; therefore, the relative time increase possibly comes from beSnake’s code initialization overhead and not the actual routing process. On the other hand, Cuccaro Multiplier and QFT, two highly connected circuits [1], obtained more than 50% rGO and rDO improvement with more than 2 and 3 times faster routing time.
@@
Focusing on ”beSnake full” columns, we see the biggest improvement in rGO, and not in rDO compared to ”beSnake fair”, and in a few algorithms, this was achieved with less routing time. This indicates that beSnake can handle highly parallelized circuits faster than less parallelized ones despite imposing more demanding routing tasks, and it can always do so with fewer additional gates. This shows that, on average, holding in place qubits to satisfy two-qubit gates, as seen in the third example in Section IV.3, does not negatively impact the gate overhead. This can be attributed to increased available space for larger topologies, which offers more alternative shortest paths around the fixed positions. Another contributing factor to this gate overhead benefit is the SWAP replacement option. Regardless, ”beSnake full” can reduce both overheads more compared to ”beSnake fair” for up to 1000 qubits while taking less time in some algorithms.

### v2

@@ V Simulations and Evaluations
With the following simulations, our goal is to test beSnake’s different functionalities and create valuable performance insights under various use cases. More specifically, we test beSnake’s capabilities with different levels of shortest path optimizations and SWAP replacement options, as well as its behavior on more connected topologies and ranges of qubit densities. Based on these insights, we move on to thoroughly compare it with the SBS algorithm on synthetic and real quantum circuits with up to 1000 qubits.
@@ V.1 Experimental Setup
The beSnake and SBS routing algorithms have been incorporated into the SpinQ compilation framework [1] and executed single-threaded on a 2.4GHz and 768GB memory server running Python 3.6.8.
Regarding the metrics used to determine their performance, we have used the total computation time of routing for each circuit, which we simply named routing time. In the case of routing time comparisons, we provide the ratio (rTime) of the two results as: (resultsa / resultsb). It should be noted that routing times can vary depending on the server’s total load and dynamic job scheduling at the time of submission. Acknowledging that, we have taken measures to maintain good time consistency between our simulations by submitting jobs in dedicated resources with no influence from other jobs.
As for overhead metrics, we have used the same gate and depth overhead definitions as in [1]. Gate overhead is calculated as the percentage relation of additional gates incorporated by the router divided by the number of gates (after decomposition). Depth overhead is expressed as the percentage of extra depth generated by the router in relation to the initial ideal circuit depth (after decomposition). Whenever there are overhead comparisons, we use a relative performance metric. Specifically, the relative gate ovehead (rGO) and relative depth overhead (rDO) is calculated as: (resultsa resultsb) / resultsa.
@@
Finally, both algorithms are benchmarked with randomly generated circuits as well as eleven well-known quantum algorithms. Regarding the randomly generated quantum circuits, they consist of two-qubit and shuttle-based Z gates in three different ratios (25, 50, and 75 two-qubit gate percent), and are sampled ten times for each data point. With this choice of random algorithms, which contain many combinations of the two gate types parallelized in different ways, we aim to stress-test both routing algorithms. As for the number of gates, we fixed it at 3000 gates since its impact on the gate and depth overhead is negligible, according to [1]. The eleven real quantum algorithms were scaled up to 1000 qubits to observe their performance in Section V.3.2. These algorithms were taken from the Qlib [55], Revlib [56], MQT Bench [57] and qbench [58] libraries.
@@ V.2 Benchmarking beSnake
This section focuses on testing beSnake’s various functionalities. Specifically, we used four time limits dictating the shortest path optimization level and compared them with taking only one shortest path in Section V.2.1. In Section V.2.2, we test the behavior of the SWAP replacement option and observe the scaling on three SWAP fidelities while maintaining the shuttle fidelity constant. Then, in Section V.2.3, we are investigating beSnake’s time performance on more connected topologies, whereas, in Section V.2.4, we are fixing the topology sizes but vary the qubit density.
@@ V.2.1 Adjusting the shortest path optimization
In the following simulations, we explore the effect various time limits have on gate overhead, as outlined in Section IV.5. Therefore, in Fig. 7, we compared the gate overhead (in blue) and total routing time (in red) for , , , and seconds time limits imposed to NetworkX.all_shortest_paths() function against just taking one shortest path with NetworkX.shortest_path(). The random circuits’ gates have been fully (ideally) parallelized based only on their dependences, and beSnake is configured without the ability of SWAP replacements (see Section IV.5), such that no other factor can influence the results. The underlying question to answer with this simulation is:
[backgroundcolor=lightgreen, linecolor=black, linewidth=1pt] Which time limit offers the best relation between gate overhead and routing time, and how does it compare with just taking one shortest path?
In Figures 7(a), 7(b) and 7(c), we can observe the gate overhead for 25, 50, and 75 two-qubit gate percentages, respectively. We observe that in all cases, the increased time limits yield a benefit in terms of gate overhead compared to taking only one shortest path, but the benefits are negligent among the time limits. As expected, higher two-qubit gate percentages result in higher gate overhead and routing time. Thus, the gate overhead gap between the one shortest path and the time limit simulations is more prominent in higher percentages. However, the overall routing time is severely increasing for each time limit. In fact, finding only one shortest path appears polynomial in time over the number of qubits with a small rGO increase of , , and on average for 25, 50, and 75 two-qubit gate percentage, respectively. To solidify this point, in Fig. 7(d), we isolated the one shortest path option, clearly showing a nearly linear behavior. A 0.05-second time limit is thus sufficient; increasing it beyond this limit does not yield significant gate overhead benefits, and further increases negatively impact the total routing time.
@@ V.2.2 SWAP replacement
In this simulation, the behavior of the SWAP replacement option on all three metrics is investigated. To test the effect of different fidelity scales, we have fixed the shuttle to , but tested SWAP gates with , , and fidelities. As explained in Section IV.5, the SWAP replacement is triggered when the accumulated shuttle fidelity is lower than the SWAP fidelity. Intuitively, we expect replacements to be triggered more often with higher SWAP fidelities. Finally, the gates in the random circuits are serialized in order to pronounce the effect in both overheads (when replacing multiple shuttles at once), and the shortest path optimization is fixed with a 0.05-second time limit. The underlying question to answer with this simulation is:
[backgroundcolor=lightgreen, linecolor=black, linewidth=1pt] Is the potential increase in routing time with beSnake worth it for architectures that support SWAP gates?
In Fig. 8, SWAP replacements improved all overheads for every two-qubit gate percentage. As expected, the SWAP replacement was utilized more often with higher SWAP fidelities and higher two-qubit gate percentages. However, there is no significant overhead difference shown in the figures for higher than (see the clustered lines), potentially due to the maximum accumulated shuttle fidelity being somewhere between and . In particular, three shuttles are needed to obtain an accumulated fidelity between and . This indicates that the BFS exploration never reached more than three layers, which is equivalent to three consecutive shuttles (similar to the example in Fig. 4(b)). Unexpectedly, we observe that the routing time has not worsened but mostly improved for higher than fidelity, possibly due to less time taken to update the intermediate representation data structure of the circuit with just one gate (a SWAP) instead of multiple shuttles. Therefore, the SWAP replacement functionality can be used without an execution time penalty in most cases. It should be noted that, given a higher qubit density on the topology, there can be more consecutive shuttles at one time, which will lower the trigger requirement and allow for even lower SWAP fidelities.
@@ V.2.3 Behavior on more connected topologies
In the following experiments, we are interested in seeing how beSnake performs on topologies with higher connectivity. Here, the comparison is made between a squared grid topology (i.e., each qubit is coupled with four neighboring qubits and with two at the corners) and a squared grid topology in which diagonal edges in every direction have been added (i.e., each qubit is coupled with eight neighboring qubits, and with three at the corners). Of course, it is expected that there will be an improvement in both overheads on the topologies with more connections. However, the increased graph size and solution space might negatively impact beSnake’s speed in exploring solutions. The random circuits are given the same way as in Section V.2.1 as well as the configuration of beSnake but only with one time limit of 0.05 seconds. The underlying question to answer with this simulation is:
[backgroundcolor=lightgreen, linecolor=black, linewidth=1pt] Does a highly connected topology adversely affect the efficiency of beSnake in quickly finding solutions?
As can clearly be seen in Fig. 9, beSnake can obtain significantly better gate and depth overhead as was expected, and surprisingly, it can also achieve it in less time. In particular, beSnake is, on average, , , and faster (rTime) for 25, 50, and 75 two-qubit gate percentages, respectively. The benefit is more pronounced with higher two-qubit gate percentages, which is expected. One explanation for the faster times can be attributed to the shorter traversed paths, hence fewer steps to take and fewer BFS exploration layers.
@@ V.2.4 Qubit density
Unlike the other simulations using a checkerboard physical initial placement pattern, we will try a different placement to vary the qubit densities here. The topology sizes are fixed, and qubits are placed next to each other in a left-to-right bottom-up approach until a particular density is reached. As for the executed circuits and beSnake’s configuration, they are similar to Section V.2.1 with a 0.05-second time limit. The underlying question to answer with this simulation is:
[backgroundcolor=lightgreen, linecolor=black, linewidth=1pt] Up to which qubit density does beSnake perform in a scalable manner based on all three metrics?
In Fig. 10, we have simulated seven levels of densities in each subfigure for four different topology sizes: a 25, 49, 64, and 100 site topology. On the y axes, we can observe the normalized mean values between 0.1 and 1 for all three metrics. Focusing on the 25-site topology, routing time (in green) increases exponentially after density, which is expected primarily due to the additional time needed to finish each BFS exploration. This becomes more severe at density, where there are only empty locations on the topology. Observing the routing time for the larger topology sizes, we see a steeper upward trend after . However, the routing time behavior until appears linear for all sizes. Notably, it becomes flatter for larger sizes, indicating that beSnake can perform better on larger topologies. This shows that beSnake’s qubit density sweet spot is around .
It should be noted, however, that for another experimental setup in regards to the hardware used to run beSnake, we might obtain a different trend and possibly an improved one if beSnake becomes multithreaded, for instance. Additionally, changing device characteristics can change beSnake’s performance. For example, increasing the connectivity will favor the run time based on the simulations in Section V.2.3. Having said that, the total run time might be reduced equivalently for all densities, and therefore, the same behavior at could still appear. Whether this sweet spot (or around that region) is universally found for any given setup is likely but left open for future work.
Focusing now on both overheads, we observe overall that beSnake favors gate overhead (in blue) over depth overhead (in orange), which is a direct outcome of the core routing strategy described in Section IV. Then, depth overhead follows a linear trend for all sizes. Surprisingly, gate overhead creates a slight curve gravitating towards qubit density. Therefore, this reinforces the previous conclusion that beSnake can operate well for more than qubit density and particularly for our particular setup. Finally, the simulations show that in terms of gate and depth circuit overhead, there are no exponential penalties for higher than densities.
@@ V.3 Comparison between beSnake and SBS
In this section, we will compare the performance of both routers for random and real circuits for up to 1000 qubits. To facilitate a fair comparison (later named as beSnake fair), the circuits are given with their gates as parallelized as possible based on SBS’s maximum capability [1]. The SWAP replacement of beSnake has been deactivated as well because SBS does not support it. Additionally, beSnake’s shortest-path optimizations are disabled, and we only use one shortest path (see Section V.2.1) to obtain the quickest routing times at the expense of overhead (see Section V.2.1). As noted before, beSnake can handle more complex routing tasks with ideally scheduled circuits based purely on their gate’s dependencies, and, for comparison purposes, we will provide such results in Section V.3.2, named as beSnake full.
@@ V.3.1 Randomly generated circuits with up to 1000 qubits
[backgroundcolor=lightgreen, linecolor=black, linewidth=1pt] Which routing algorithm will more likely produce lower gate and circuit depth overhead for scalable architectures that go up to 1000 qubits, and at what relative time cost?
In Figures 11(a) and 11(b), the gate overhead and depth overhead of both algorithms (marked as SBS and beSnake) are shown for the different two-qubit gate percentages (in yellow, blue, pink). We can observe beSnake (circle markers) a relative improvement in gate overhead (rGO) of and in depth overhead (rDO) of on average compared to SBS (square markers). Additionally, we observe beSnake’s relative benefit in both overheads increasing with higher qubit counts. Looking at the three two-qubit gate percentages, we see an increased gate overhead improvement with higher percentages (see the distance between curves of the same color). More specifically, the average rGO is at % for 25%, for 50%, and for 75%, and the average rDO is calculated at for 25%, for 50%, and % for 75%. beSnake will only be, on average, slower (rTime) than SBS at 25%, at 50%, and at 75%. Notably, both algorithms’ routing time starts getting closer around and after 900 qubits, indicating that beSnake takes almost the same amount of time at 1000 qubits, especially for more complex routing tasks (i.e., 75 two-qubit gate percentage shown in Fig. 11(c)). Considering the performance gains, beSnake is a more attractive solution for large-scale architectures overall. Finally, as discussed before, in these comparison simulations, beSnake is not utilized fully; hence, we expect even better performance.
@@ V.3.2 Real algorithms with up to 1000 qubits
Ratio: (Average of two-qubit gates + Average of Z gates) / Average of two-qubit gates. beSnake fair: beSnake is configured such that it matches the maximum capabilities of the shuttle-based SWAP (SBS) algorithm. beSnake full: beSnake is used at its full capacity. rGO [%] and rDO [%] refer to relative gate and depth overhead, respectively: (SBS average results beSnake average results) / SBS average results. rTime: SBS average time / beSnake average time.
@@
In the following simulations, we compare the performance of SBS over beSnake for real algorithms with up to 1000 qubits. Previously, the random circuit set was produced to stress test the capabilities of both routing algorithms; however, the results are not representative of routing for real algorithms, which might exhibit different circuit characteristics. Similarly to Section V.3.1, we will compare both routers fairly to see their relative performance. However, we will also assume ideally parallelized circuits and enable the SWAP replacement option (at 99.95% SWAP fidelity) in order to test beSnake at its full capacity. The underlying question to answer with this simulation is:
[backgroundcolor=lightgreen, linecolor=black, linewidth=1pt] How much better can beSnake perform over SBS on real quantum algorithms, and is it worth the potentially increased routing time?
@@
Table 2 summarizes the relative performance of beSnake over SBS (used as a baseline) in two configurations, one fair comparison, beSnake fair, and one scenario with beSnake at its full capacity, beSnake full. Due to algorithm availability, we have specified each algorithm’s qubit range and incremented qubit steps on the second column. One the next 3 columns, we have provided the average two-qubit and Z gates together with their ratio. However, it should be noted that the number or percentage of gates is not the only performance indicator or predictor. To get a full picture of their features, a deeper analysis has to be conducted with the help of the algorithm’s internal interaction and dependency graph [1, 58]. In this simulation, we are focusing on the performance comparison of the two routers rather than on specific quantum algorithm performance or how these algorithms scale.
Moving on to the rest of the columns, we provide the relative performance of beSnake over SBS for gate overhead, depth overhead, and routing time as defined in Section V.1. For example, beSnake obtained a 16% gate overhead and 29% depth overhead improvement over SBS under a fair comparison for Grover’s algorithm. As for routing time, beSnake is 42% slower (rTime = 0.58) than SBS.
@@
Looking only at the fair comparisons overall, we observe an improvement in both rGO and rDO across all algorithms, with up to a 67% improvement in rGO and up to 45% improvement in rDO. On the one hand, rTime results for Deutsch-Josza (up to 1000 qubits) and for Bernstein-Vazirani (up to 50 qubits) seem relatively high for their low average two-qubit and Z gate values. Evidently, these algorithms do not require a lot of routing; therefore, the relative time increase possibly comes from beSnake’s code initialization overhead (i.e., loading libraries, allocating memory, initializing variables, and setting up data structures) and not the actual routing process. On the other hand, Cuccaro Multiplier and QFT, two highly connected circuits [1], obtained more than 50% rGO and rDO improvement with more than 2 and 3,5 times faster routing time.
@@
Focusing on beSnake full columns, we see the biggest improvement in rGO, and not in rDO, compared to beSnake fair, and in a few algorithms, this was achieved with less routing time. This indicates that beSnake can handle highly parallelized circuits faster than less parallelized ones despite imposing more demanding routing tasks, and it can always do so with less gate overhead. This shows that, on average, holding in place qubits to satisfy two-qubit gates, as seen in the third example in Section IV.3, does not negatively impact the gate overhead. This can be attributed to increased available space in larger topologies, which offers more alternative shortest paths around the fixed positions. Another contributing factor to this gate overhead benefit is the SWAP replacement option. Regardless, beSnake full can reduce both overheads more compared to beSnake fair for up to 1000 qubits while taking less time in some real algorithms.